OpenStudy (anonymous):

A hydrate of Na2CO3 has a mass of 4.31 g before heating. After heating, the mass of the anhydrous compound is found to be 3.22 g. Determine the formula of the hydrate and then write out the name of the hydrate.

OpenStudy (anonymous):

I found the mass of the water, it's 1.09 g, but I am lost from there on.

OpenStudy (aaronq):

find how many moles there are of the anhydrous compound and moles of water, separately. Compare them

OpenStudy (anonymous):

So I would start with 22.989g/3.22g, and then multiple the that answer by 2, to get the moles of the Na2? Or would I multiple 22.989 by 2 and then divide?

OpenStudy (anonymous):

Or does the 2 not matter?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

i meant find the moles of \(Na_2CO_3\) and the moles of \(H_2O\), not of the individual atoms

OpenStudy (anonymous):

I'm sorry, moles have always confused me a bit. How would I start with finding the moles of Na2Co3?

OpenStudy (aaronq):

you need the molar mass (which is the sum of the individual molar masses of the atoms in their proportions, so like 2 times the molar mass of Na + molar mass of C + 3 times molar mass of O). use this formula: \(\sf moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (anonymous):

Okay so molar mass of\[Na _{2}CO _{3}\] is 105.985 g?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

and then of H2O it would be 18.0148

OpenStudy (anonymous):

Thank you for being patient with me aha

OpenStudy (aaronq):

yes, you can just use 18 g/mol and no problem! haha

OpenStudy (anonymous):

So from here, where do I go ?

OpenStudy (anonymous):

Divide the mass of original compound by the mass?

OpenStudy (aaronq):

divide the mass of the anhydrous compound by the molar mass of Na2CO3 divide the mass of the water by it's molar mass

OpenStudy (anonymous):

3.22/ 106g = .030 g 1.06/18g= .060 g

OpenStudy (anonymous):

Divide by the smallest number in the ratio to get a whole number ratio , right?

OpenStudy (aaronq):

good stuff! remember that this is moles not grams, \(\sf \dfrac{3.22~g}{ 106~g/mol}\) = .030 mol \(\sf \dfrac{1.06~g}{18 ~g/mol}\)= .060 mol

OpenStudy (aaronq):

yeah

OpenStudy (anonymous):

Oh yeah right , thanks! so from here, it's .030 mol/ .030 mol = 1 mol .060 mol/.030 mol = 2 mol

OpenStudy (anonymous):

Na2CO3 * 2H2O or Sodium carbonate dioxide?

OpenStudy (anonymous):

*dihydrate

OpenStudy (aaronq):

haha yes i was gonna say thats wrong but you corrected yourself. good job

OpenStudy (anonymous):

Awesome! This problem killed me but we finally finished it. Thanks so much for your help!!!

OpenStudy (aaronq):

no problem !